what are three consecutive positive numbers such that the product of the first and third, minus the second,is one more than six times the third
Answer by CubeyThePenguin(3113)

consecutive numbers: (x-1), x, (x+1)

(x-1)(x+1) - x = 1 + 6(x+1)
x^2 - x - 1 = 6x + 7
x^2 - 7x - 8 = 0
(x - 8)(x + 1) = 0

x is positive, so the numbers are 7, 8, and 9.